Fermilab Shutdown Report
7/24/09 – 7/31/09
LINAC
Shutdown Week 7 – On Schedule
- Preaccelerator
- Recovered the H- Source from the power outage
- H - is operational putting out 54 mA
- Finished I- Source clean up/assembly
- Painting of the I - & elevator pit
- Low Energy
- RF power turned on in the LRF1-4 & Buncher
- Completed Linac/ Neutron Therapy Facility (NTF) beam sign off
- Successfully accelerated beam to NTF
- Carefully initiated the RF power turn on for LRF5
- High Energy
- Completed work on Klystron RF station 6
- Initiated the HE QPS work
- Ongoing work cleaning the Klystrons heat exchangers
Booster
Shutdown Weeks 7 – On Schedule
- Tunnel Activities
- Correction element magnet work
- Vacuum pump down is progressing, although some areas are slower to improve
- All new correctors on 473 modules
- Water connections complete
- Database updates continue
- GMPS work continues 85% complete
- East Gallery LCW line hydro-test
- Booster 95 LCW circulating and cleaning
- Clean up crew cleaned Booster tunnel
- Tunnel sump discharge line repaired
- Next Week:
- Corrector work continues
- Vacuum pump down, leak checking
- Controls work
- Checking magnet fields
- Finish connecting cables to power supplies
- Power supply testing
- GMPS work continues
- Repair leaks in East gallery LCW system
- CUB Booster LCW system, operational
Antiproton Source
Accumulator 4-8GHz DP Kicker Upgrade: ~75% complete
After the A50-A60 bake was completed, tests on both kicker tanks indicated internal electrical problems. Vacuum was let up in A50 and the tanks disassembled for inspection. We found that the microstrip leads from the vacuum feed throughs to the arrays inside the tanks failed during the bake. Repairs are underway. A50 will be baked again next week.
Debuncher Septum Magnet Replacement: ~85% complete
The repaired magnet has been installed and surveyed. As soon as the welding is completed and vacuum is restored, pulse testing will begin.
IB7 (AP2 dipole magnet) Replacement: ~99% complete
The magnet is in and surveyed. All that remains to do is pump down the line (in progress), flush the magnet, and reconnect the LCW.
Other Work
- RF system work. Maintenance work on various RF systems continues.
- LCW leaks soon to be addressed.
- Target SEM work continues.
- Following our power outage last weekend, Debuncher vacuum in the 30 sector had a very hard time recovering. We will be looking into this soon.
- Cryostat vacuum O-ring replacement work to begin soon. This must be completed before A60 LN2 can be reconnected.
- Pbar LCW pumps were not turned back on after last weekend’s power outage until we inquired about it on Tuesday. Although this did not cause a problem, it could have. We still don’t know why it took so long … ?
- Although our kicker array problem cost us ~1.5 weeks, we are still on schedule to finish before the end of the shutdown!
Main Injector
Tunnel Work MI-8
- Multiwires
- 830, 839 & 840 installed
- 810, 811, 813 & 814 ready for installation on Monday, Aug. 3 rd
- We will be replacing two Texas wires that are currently being referenced and replacing the 851 multiwire
MI-40
- Pipe fitters are fabricating and installing cooling systems.
- Pipe arrived in less that ideal condition requiring extra prep/cleaning.
- First kicker delivery expected Monday
- We estimate 6 kickers delivered with time to install. 7 th kicker may be to late.
MI eCloud (MI-52)
- Lambertson is continuing to bake
- All eCloud components are ready for installation
- Many problems with ferrite absorbers and cloud detectors
Other Work
- MI-30 collimator LVDT investigation continues
- MI-30 collimator mask installation scheduled to begin Aug. 24 th
- MIPS & MIRF work going well, on schedule.
- Lead blankets: remove from MI-40 on Aug. 10 th
- Lead blankets: remove from MI-10 on Aug. 13 th
- After lead is removed LCW re-hosing will finish at those locations.
- LCW fill set to begin Aug. 18 th and conditioning on Aug. 21 st
Recycler
Recycler
- Pot RR129 (i.e. MI-10 injection kicker) (80% complete)
- Potted flange is curing
- Leak check new (larger) flourinert vials today
- Move 6 magnets (complete)
- Vacuum maintenance
- Leak checking 800 line
- One leak was found… but disappeared. Applied VAC seal just in case.
- Next: Leak check 100-103 (stochastic cooling kickers)
- Power supplies maintenance
- CIA crate activities
- HLS installation at MI-10
- ALARA plan has been reviewed
- Installation within the next two weeks (1-2 days)
Electron Cooling
- Pelletron maintenance (60% complete)
- Encountered problem with the desiccant regeneration controls system yesterday
- New filters were installed on the transfer and cooling skids
- Decided on installing a new 2” GVM as a ‘hot spare’ instead of the repaired 4” GVM
- Water/Glycol system cleanup and upgrade ( Complete)
- BUT… failed cooling test
- More troubleshooting needed; likely just trapped air
- NMR probes replacement (90% complete)
- One last probe needs to be tuned in
- Ecool is on track for starting up on August 19 th (gas transfer)
- Interlocks tested Tuesday (U-bend mode only)
Recycler Pump Out
- Vacuum pump out started on Monday 7/27/09 at 401-409
Tevatron

Other Items
- Start replacement of 8 bad stands expect 1 per day
- Un Roll TeV magnets in A-1 next week
- Finishing in F1 all test OK fixing last insulating vacuum leak
- Hipot early next week
- Vertical crystal collimator bake out nearly complete
- Repairing HLS systems around Tevatron D0 system up B0 partially, complete need to have IMU closed to finish
- Will have day and night shift for vacuum work started
- Still to do Dump resistor at F46
- Install vertical xtal collimator soon
NuMI
Shutdown Week 7
- Remove all R-blocks from Target Pile for survey work
- Complete survey of Horn 2, Horn 1, and the Target
- Complete additional survey work of Pre-Target Area
- Replace supply water lines on Horn 1
- Remove water lines on Target
- Replace motor and desiccant wheel on Decay pipe Munters Unit #1
- Start repair of pump on MI65 Evaporator Mixing Tank
- Completed PMs on MI65 & MINOS Elevators including roller replacement
- Continue Installation of AH Dehumidification System
- Controls System
- Pumps & Heat Exchanger
- Complete installation of rope filter system for ground water cooling leg to Absorber Hall
- EPB Dipoles PEEK Installation Continues
- HV101-3 is out & PEEK replacement in process
- Shielding off of HV101-4 & HV101-5
- Completed minor adjustment to shielding 101-6
- Survey completed fiducialization of new Texas Multi Wires for PM114 & PMTGT & start pre-installation checkout
- Completed installation of Hadron Monitor Fixture and complete pre-operational testing
Cryogenic System
TeV at LN2 Temperatures

Tevatron Operations and Maintenance Status
- Continued support AD/MSD for vacuum repairs: D3 is back at LN2 Temp. A3 is ready for cooldown. E3 and F1 will soon be transferred from MSD to Cryo for restoration to LN2. A1 is transferred to MSD; A2 is warming up
- LN2 inventory control and controlling magnet temperatures at LN2
- Expanders and Instrumentation group: All Tevatron wet & dry expanders complete; all planned flywheels replaced. CDF engines are complete. DZero engines will be done next week. Still need to sting all U-tubes. Still need to complete transducer maintenance. Enormous backlog of shop repair and overhaul work
- Electrical group: Continuing upgrade of Cutler Hummer 120V power, 2 compressors left. Continuing maintenance of cryo I/O crate, electrical equipment cabinets, various low voltage electrical/electronic maintenance
- Compressor group: Mycom Compressor relief valves replaced (3/8). All IHI cold compressors replaced; 2/5 IHI has been overhauled; C4 IHI is installed to replace reciprocating compressor
- Ongoing numerous FRIG/Compressor controls/instrumentation maintenance/repairs
---------------------------------------------------------------------------------------
- Pending RT houses and relief maintenance, Tevatron cryo is on schedule
CHL HELIUM and NITROGEN PLANTS
- 90% of all scheduled routine mechanical and electrical maintenance is complete
Non-complete critical items
- Scheduled recycle TSI 2.5MW compressor overhaul for Aug 10 by outside subcontractor
- Expect delivery of repaired turbo expanders for the CHL coldbox-2 on Aug 17. Following by installing repaired turbines into coldbox-2. Following by test run and commissioning coldbox-2 into service. Backup plan – commission coldbox-1 into service
- Scheduled repairs of compressor D for Aug 10 by outside subcontractor
- Continuing maintenance of degraded cooling tower for the nitrogen plant. Followed by testing and commissioning
- Installation of the new plate and frame ICW heat exchanger
- Repairs of the Nitrogen plant Marley water tower
- Scheduled maintenance and commissioning of the helium pump
- Pending successful TSI and coldbox-2 repairs, maintenance, CHL cryo is on schedule
Controls
Completed This Week:
- Control System weekend outage
- Control system recovery - Monday
- Linac startup support
- General Controls Maintenance
Next Week:
- TLG S/W Testing on TLGTEST – MOOC 4.0 update
- Preparation for TLG boot Monday 8/10
- Network Isolation Test – Thursday 8/6
- 30 min. no email, internet access from AD networks
- Support Linac startup activities
- General Maintenance
MTA
The permanent magnet has been built and is on a stand waiting to be measured and painted. It maybe installed near the end of next week.
FESS
Electrical and High Voltage
- The Kautz Road buss duct cleaning is ongoing this week. Small oil leaks were found. This is “Siemens’ Warranty “work
- Re-lamping is complete at MI TEV & Pbar some repairs, Work in Progress (WIP). Switchyard re-lamp, WIP
- Feeder 45 will be down on Saturday (8/1/09) for transformers and switch cleaning –TeV Off for 10 hrs, CO ASSY off 8hrs
Mechanical
- TEV Ring Service Bldg. Pond Pump and Cryo pump repairs, WIP
HVAC
- •TeV Ring Service Bldg. HVAC and RTF De-humidifier PM, WIP
Engineering
- The 12” ICW Crossover Connection has been restored at RD. A and B
- New ICW Break at TPL, Fire Protection and Cooling down to HIL, PW8, Site 49, NEB, NS4, PS5 - WIP
- Kautz Road 345Kv line dampers, WIP. Static line support damaged
CUB
- LCW pump PMs, WIP
- Linac LCW Chiller status – Tube & tube sheet replacement recommended, 4 weeks to replace. CUB reconfiguration will provide Linac 55 Degree LCW via back-up heat exchanger
- Chiller Cleaning and Eddy Current Testing on schedule
- 95 Booster and 95 P Bar LCW Heat Exchanger Cleaning and Inspection on schedule
Casey’s Pond/ICW System
- Isolated FCC from ICW Grid and supplied ICW from C4. Supplied WH 8 th Floor Computer Room from WH basement. Chiller on Domestic Water. Shut down ICW System, cleaned pump wet well, removed approx 10,000 lbs of Zebra Mussels, mud and debris
Switchyard 120 – Work has not started
Collider Detector at Fermilab
- The end plugs will be installed around 8/12/09
D-Zero – On Schedule
Mechanical – On Schedule
Instrumentation– On Schedule
Alignment

Alignment
- As found shielding MI10 and MI40 - Working on today
- Survey of the berm for radiation shielding at MI10 after the civil construction is complete.
- Survey of the berm for radiation shielding at MI40 after the civil construction is complete
- Align 7 new kicker magnets to be installed in the MI40 area. See MI personnel for details. 7 Components
- After the shutdown is complete we may need to move magnets to adjust our orbits in the Main Injector and Recycler. This could be as much as 5 magnets in each machine (10 total) Shutdown 2009
- As-found all the Main Injector and Recycler components in cells102-109. The essential components are: Recycler = magnets and cooling tanks, no trims. 40 Components
- As-found all the Main Injector and Recycler components after construction back fill in cells 341-404. Align all kickers and new beam pipe. 40 Components
- Realign 6 EPB dipoles @ (HV101 magnet string). 6 Components
- Align 3 replacement Fermi multiwires at PM101, PM105, and PM114. 3 Components
- Align a new Texas multiwire at PMTGT. 1 Component
- Re-align Fermi multiwires at PM121 after stand adjustment (and re-check adjacent instrumentation). 1 Component
- Layout the location of 6-7 magnets in the NuMI Stub for LBNE ( DUSEL) beamline magnet positioning optimization (estimated total time 1day). 7 Components
- Stickmic survey of Muon system including CF and EFs after closing detector. The start and finish dates are preliminary and depend on shutdown progress. Shutdown 2009
- As found 2 EPB dipoles in enclosure B (VH94) before reworked EPD dipoles. Align 2 new dipoles
- Align new permanent magnet for MTA to be installed on the Linac side. See MTA personnel for details.
Safety
Safety
- Linac and NTF approved for operations
- Accelerator restart sheets are available from ES&H
Interlock Work
- Single Line Electrical Drawing Verifications
- P-Bar 95% complete
- Tevatron not started
- Cables for future MuCool Rep-Rate Limiter interlock installed
- Interlocks for the Booster laser profile monitor completed
- Misc. Tevatron Interlock repairs
- Tevatron ground fault diagnostics installation
- Safety System Testing
- Pelletron testing for U-Bend operations completed
- A0 Photo injector and North Cave testing completed
- MDB 325MHz RF and Horizontal Test Cave testing completed
Shutdown Safety
Continue using the Take Five principles before, during, and after work activities to:
- Avoid Injuries
- Reduce Environmental Impacts
- http://www.fnal.gov/takefive/
|